# Implicit Dynamical Flow Fusion (IDFF) for Generative Modeling

### v1

@@ 3 Implicit dynamical flow fusion model (IDFF)
Increasing the sampling steps to improve generative efficiency in FMs while preserving the target distribution’s accuracy is crucial. To achieve this, we introduce a learnable Hamiltonian momentum term, denoted as $\bm{\xi}_{t}$, into the vector field $\hat{{\mathbf{v}}}_{t}$ and define a new vector field for FMs as:
$\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})=\sqrt{1-\sigma^{2}_{t}}\hat{{\mathbf{v}}}_{t}({\mathbf{x}})+\sigma^{2}_{t}\bm{\xi}_{t},$ (6)
where the marginal vector field $\hat{{\mathbf{v}}}_{t}({\mathbf{x}})$, associated with the conditional vector field $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})$, generates the marginal probability path $p_{t}({\mathbf{x}})$ of these conditionals and $\sigma_{t}=\sigma_{0}\sqrt{t(1-t)}$. We define the momentum term $\bm{\xi}_{t}=\gamma\nabla_{x}\log p_{t}({\mathbf{x}}_{t})$ to guide the vector field towards the gradients of the log probability of samples ${\mathbf{x}}_{t}$, using a weight factor $\gamma$, as depicted in Figure 2.
When $\sigma_{t}$ is set sufficiently close to 0 for $t=\{0,1\}$, $\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})$ converges to the standard $\hat{{\mathbf{v}}}_{t}({\mathbf{x}})$, ensuring the same marginal distributions at $t=0$ and $t=1$ as in CFMs. As demonstrated here, incorporating such an auxiliary variable into Langevin sampling methods aligns with well-established HMC simulated sampling methods. This strategy reduces the steps required to achieve a high-fidelity distribution Hoffman et al. (2014); Neal (2012).
By substituting the new conditional vector field $\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})$, defined in equation 6, in the continuity equation as described in Appendix A.1.1, we introduce the vector field ${\mathbf{w}}_{t}$, which defines the IDFF (Implicit Dynamical Flow Fusion) probability flow ODE as:
${\mathbf{w}}_{t}=\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\mathbf{x}}_{t})=\sqrt{1-\sigma_{t}^{2}}\hat{{\mathbf{v}}}_{t}({\mathbf{x}})+\frac{(2\gamma-1)}{2}\sigma_{t}^{2}\nabla\log p_{t}({\mathbf{x}}_{t}),$ (7)
where the conditional process starts with the initial condition ${\mathbf{x}}_{0}\sim\mathcal{N}(0,\textbf{I})$ and follows a specific conditional probability path $p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$ to ultimately generate ${\mathbf{x}}_{1}$.
@@
Given that $\hat{{\mathbf{v}}}_{t}({\mathbf{x}})$ represents the vector field that generates $p_{t}({\mathbf{x}})$ and $\sigma_{t}$ is set sufficiently close to 0 for $t=\{0,1\}$. The IDFF probability flow ODE defined in equation 7 follows the same marginal distribution as the CFMs with the ODE defined in equation 3.
3.1 Training Objective for IDFF
As the CFM literature outlines, learning a conditional vector field $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$ is faster and preserves the same marginals. Consequently, we can define the conditional probability OT path $p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$ to act as a probability bridge between a source point ${\mathbf{x}}_{0}$ and a target point ${\mathbf{x}}_{1}$ (Tong et al., 2023; Lipman et al., 2022).
An efficient approach to selecting the probability path $p_{t}(.)$ from a displacement perspective is to choose the path with minimal displacement between $p_{0}(.)$ and $p_{1}(.)$ by defining it as a stochastic interpolator between ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{0}$, given by:
$p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})=\mathcal{N}({\mathbf{x}}\mid t{\mathbf{x}}_{1}+(1-t){\mathbf{x}}_{0},\sigma_{t}^{2}\textbf{I}),\ \ \sigma_{t}=\sigma_{0}\sqrt{t(1-t)}$ (8)
$\hat{{\mathbf{v}}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1})=\frac{{\mathbf{x}}_{1}-{\mathbf{x}}}{1-t}$ (9)
This definition aligns with the conditional OT paths between ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{0}$, which have an equivalent conditional probability flow to the marginal $p_{t}(.)$ of flow $\hat{{\mathbf{v}}}_{t}(.)$ (Pooladian et al., 2023).
@@ 3.1 Training Objective for IDFF
For large datasets, computing this OT map can be challenging (Cuturi, 2013; Tong et al., 2020). A minibatch approximation of OT, similar to (Fatras et al., 2021), can address this. While minibatch OT introduces some errors compared to the exact OT solution, it has proven effective in numerous applications such as domain adaptation and generative modeling (Genevay et al., 2018). Therefore, we use minibatch OT to approximate the map between ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{1}$.
@@
Current CFM objectives utilize the loss function defined in equation 5, which optimizes the prediction of the vector field. Instead of directly learning $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}\mid{\mathbf{x}}_{1};\theta)$ using equation 5, we propose defining a denoising neural network $\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t;\theta)$ and reparameterizing $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}\mid{\mathbf{x}}_{1};\theta)$ as follows:
$\hat{{\mathbf{v}}}_{t}({\mathbf{x}}\mid{\mathbf{x}}_{1};\theta)=\frac{\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t;\theta)-{\mathbf{x}}}{1-t}$ (10)
We suggest a loss function comprising two terms: the first term directly optimizes the denoising model $\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t;\theta)$, with ${\mathbf{x}}$ as the noisy input, and the second term is a scoring model that approximates $\nabla_{x}\log p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$. Optimizing $\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t;\theta)$ and $\epsilon({\mathbf{x}},t;\theta)$ directly implies optimizing the vector field $\tilde{{\mathbf{v}}}_{t}({\mathbf{x}}\mid{\mathbf{x}}_{1};\theta)$, which is the intuition behind naming the IDFF model as an implicit model.
Considering that both $\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t;\theta)$ and $p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$ are easy to sample from, we can use neural networks to approximate both and train them with the following loss function:
@@
$\mathcal{L}_{\text{IDFF}}(\theta)=\mathbb{E}_{...}\big{[}\beta(t)^{2}\|\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t;\theta)-{\mathbf{x}}_{1}\|^{2}+\lambda(t)^{2}\left\|\epsilon({\mathbf{x}},t;\theta)-\nabla_{x}\log p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})\right\|^{2}\big{]}$ (11)
@@
Where $\{t\sim{\mathcal{U}}(0,1),{\mathbf{x}}_{0}\sim\mathcal{N}(0,\textbf{I}),{\mathbf{x}}_{1}\sim p_{data},{\mathbf{x}}\sim p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})\}$. Here $\{\lambda(\cdot),\beta(\cdot)\}$ are two sets of positive weights. See Appendix A.1.2 for more details about the loss function. The training process is described in Algorithm 1.
3.2 IDFF time-series adaptation and sampling process
To use IDFF for time-series applications, we need to modify the training and sampling algorithms for IDFF. To accomplish that, we define another random variable, $k$, which is a discrete variable and can take values of $\{1,...,K\}$ with equal probability, where $K$ is the length of a time series. Therefore, $k$ represents the index of the data sample in a time series. We also have $t$ as a continuous time variable interpolated between two subsequent values of $k$ (i.e., $k$ and $k-1$). Therefore we have $\forall k\in\{1,...,K\}\rightarrow p(t\mid k)\sim{\mathcal{U}}(0,1)$. This conditioning does not change the original assumption over $t$ since $k$ can take any values up to the time-series length with the sample probability. With this in hand, we can directly pass the $k$ as additional inputs to $\hat{\bm{\epsilon}}_{1}({\mathbf{x}},t,k;\theta)$ and $\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t,k;\theta)$ and suggest Algorithm 1 for training and Algorithm 2 for sampling time-series data with IDFF. Note that for static data, the sequence length ($K$) equals one; otherwise, it equals the time series length.
For a fixed $K=1$, once we have trained $\hat{{\mathbf{x}}}_{1}(.;\theta)$ and $\bm{\epsilon}(.;\theta)$, we can generate samples simulated by an SDE from source samples ${\mathbf{x}}_{0}$. This procedure is outlined in Algorithm 2. The SDE is defined as $d{\mathbf{x}}_{t}={\mathbf{w}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1})dt+\sigma_{t}d\mathbf{w}$, with initial samples ${\mathbf{x}}_{0}$ and target samples ${\mathbf{x}}_{1}$ at $t=1$. Therefore, we can draw samples from $p({\mathbf{x}}_{0})$, run them through the SDE, and obtain the final samples ${\mathbf{x}}_{1}$.
@@ 3.2 IDFF time-series adaptation and sampling process
First, we draw ${\mathbf{x}}_{0}\sim\mathcal{N}(0,\mathbf{I})$, then compute $\hat{{\mathbf{x}}}_{1}({\mathbf{x}},0,k;\theta)$. Next, we calculate $\hat{{\mathbf{v}}}_{t}$ using equation 10, followed by the computation of ${\mathbf{w}}_{t}$. After calculating ${\mathbf{w}}_{t}$ and using the SDE, we compute the mean for the next time step as $\bm{\mu}_{t+\Delta t}=({\mathbf{x}}_{t}+{\mathbf{w}}_{t}\Delta t)$, and finally draw a sample from a small neighborhood of it: ${\mathbf{x}}_{t+\Delta t}\sim\mathcal{N}({\mathbf{x}}_{t}+{\mathbf{w}}_{t}\Delta t,\sigma_{t}^{2}\Delta t\mathbf{I})$.

### v2

@@ 3 Implicit Dynamical Flow Fusion Model (IDFF)
IDFF aims to improve sampling efficiency while enhancing sample quality of OT-CFMs through two key innovations: (1) the incorporation of a learnable Hamiltonian momentum term, $\bm{\xi}_{t}$, into the conditional vector field, which enhances sampling efficiency; and (2) the proposal of a new objective function that learns the IDFF’s vector field in input space.
3.1 Learning CFMs with Momentum
@@
To improve sampling efficiency while preserving distribution accuracy in CFMs, we introduce a learnable Hamiltonian momentum term, $\bm{\xi}_{t}$, integrated with the original vector field $\mathbf{v}_{t}(\mathbf{x}_{t})$. This integration yields a new, more flexible vector field:
$\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})=\sqrt{1-\sigma^{2}_{t}}{\mathbf{v}}_{t}(\mathbf{x}_{t})+\sigma^{2}_{t}\bm{\xi}_{t},$ (6)
where $\sigma_{t}$ is a time-dependent interpolation factor balancing ${\mathbf{v}}_{t}$ and $\bm{\xi}_{t}$.
Interpolating the vector field: We define the interpolation factor $\sigma_{t}$ as $\sigma_{t}=\sigma_{0}\sqrt{t(1-t)}$, where $\sigma_{0}$ is a hyperparameter that controls the overall influence of the momentum term. While the IDFF vector field converges to OT-CFMs in either limit, outside of it, it enables the flow to leverage the momentum variables to guide the flow based on the momentum variables. Our empirical results demonstrate that this approach significantly improves sampling efficiency by substantially reducing the NFEs required. Our approach also overcomes the limitations of uniform weighting across trajectories by prioritizing samples near $t=0,1$, allowing finer attention to complex structures; see Figure 2.A.
Designing momentum variables: The term $\bm{\xi}_{t}$ is designed to guide the new vector field toward the correct probability flow for the next time step, ensuring its role in accurately modeling the distribution. There are many potential choices to make here. Still, we take inspiration from Hamiltonian Monte Carlo, which showcases how to leverage momenta to rapidly sample complex multi-modal posterior distributions Neal (2012b). A potential design for $\bm{\xi}_{t}$ is $\bm{\xi}_{t}=\gamma\nabla_{x}\log p_{t}(\mathbf{x}_{t})$, which aligns with the gradient of the log probability of the samples $\mathbf{x}_{t}$. If $p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{1})$ is a Gaussian with variance $\sigma_{t}^{2}$, $\bm{\xi}_{t}$ can be simplified to $-\frac{\bm{\epsilon_{0}}}{\sigma_{t}}$, where $\bm{\epsilon_{0}}\sim\mathcal{N}(\bm{0},\bm{\text{I}})$. Alternatively, higher-order choices for momentum terms could capture even more complex interactions in the flow. $\bm{\xi}_{t}=\gamma\nabla_{x}\log p_{t}(\mathbf{x}_{t})$ drives the vector field toward regions of high probability, thereby improving the accuracy of the generated samples. $\gamma$ serves as a weighting factor that controls the influence of this term.
To maintain the same probability path as the original CFM, the vector field must satisfy the continuity constraint defined in equation 4. By incorporating $\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})$ into the continuity equation 4 (detailed in Appendix A.1.1), we derive the IDFF probability flow ODE:
$\mathbf{w}_{t}(\mathbf{x}_{t})=\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}(\mathbf{x}_{t})=\sqrt{1-\sigma_{t}^{2}}{\mathbf{v}}_{t}(\mathbf{x}_{t})+\frac{(2\gamma-1)}{2}\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t}),$ (7)
To ensure that our new marginal vector field $\mathbf{w}_{t}(\mathbf{x}_{t})$ converges to the standard ${\mathbf{v}}_{t}(\mathbf{x}_{t})$ at the start and end points of the process, $\sigma_{t}$ must approach zero as $t\to{0,1}$. This convergence is crucial for maintaining consistent marginal distributions at these points.
Therefore, ${\mathbf{w}}_{t}({\mathbf{x}}_{t})$ now replaces the vector field ${\mathbf{v}}({\mathbf{x}}_{t})$ in the probability flow ODE (equation 3), enabling the construction of a new generative process that follows the same probability path $p_{t}(\mathbf{x}_{t})$. This process starts with $\mathbf{x}_{0}\sim\mathcal{N}(0,\mathbf{I})$ and follows $p_{t}(\mathbf{x}_{t})$ to produce $\mathbf{x}_{1}$. The resulting model thus implicitly models the score function using flow matching which motivates the name Implicit Diffusion Flow Fusion.
@@
Given that ${\mathbf{v}}_{t}(\mathbf{x}_{t})$ generates $p_{t}(\mathbf{x}_{t})$ and $\sigma_{t}$ approaches 0 for $t\in{0,1}$, the IDFF probability flow ODE (Eq. 7) follows the same marginal distribution as the CFMs with the ODE defined in equation 3.
Proof Schema for Lemma 1.
Our goal is to show that the IDFF probability flow ODE shares the same marginal distribution as CFMs when $\sigma_{t}$ is small for $t\in{0,1}$.
  • Step 1: Define the vector fields. We introduce $\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})$ governed by the $\bm{\xi}_{t}$ and ${\mathbf{v}}_{t}(\mathbf{x}_{t})$ from the original CFM.
  • Step 2: Apply the continuity equation. By substituting $\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})$ into the continuity equation, we derive the IDFF ODE, ${\mathbf{w}}_{t}(\mathbf{x}_{t})$.
  • Step 3: Analyze boundary conditions. We show that at $t=0$ and $t=1$, where $\sigma_{t}$ approaches 0, the IDFF ODE converges to the CFM-ODE, ensuring the same marginal distributions.
Appendix A.1.1 provides the complete proof. This lemma and its proof establish the consistency of IDFF concerning existing CFMs. The approach showcases how to design a new flow augmented with momentum variables with an interpolant that guarantees that it matches OT-CFM in the limit. In what follows, we discuss how to parameterize and learn such a flow.
3.2 Learning Flows in Input Space
One of the challenges in our formalism is learning ${\mathbf{w}}_{t}$. This challenge arises because ${\mathbf{v}}_{t}$ operates in vector space while $\bm{\xi}_{t}$ remains in the input space. Rather than training one part of the model in vector space and the other in input space, we define a new objective defined entirely within the input space to address this. Learning the vector field in the input space implies optimizing the vector itself, which is the reasoning behind naming the IDFF model an implicit model of dynamical flows. We begin by defining an OT path $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$, which acts as a probability bridge between $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$. This path is a stochastic interpolator between the source $\mathbf{x}_{1}$ and target $\mathbf{x}_{0}$, as follows:
$p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})=\mathcal{N}(\mathbf{x}_{t}\mid t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}),\ \ \sigma_{t}=\sigma_{0}\sqrt{t(1-t)}$ (8)
The corresponding conditional vector field is given by:
$\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})=\frac{\mathbf{x}_{1}-\mathbf{x}_{t}}{1-t}$ (9)
@@ 3.2 Learning Flows in Input Space
This definition aligns with the conditional OT paths between $\mathbf{x}_{1}$ and $\mathbf{x}_{0}$. For large datasets, computing this OT map can be challenging. We employ a minibatch approximation of OT to address this, similar to (Fatras et al., 2021). While this introduces some errors compared to the exact OT solution, it has proven effective in numerous applications such as domain adaptation and generative modeling (Genevay et al., 2018).
@@
Current CFM objectives typically use a loss function (defined in equation 5) that optimizes the prediction of the vector field. We propose an alternative approach: defining a denoising neural network $\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)$ and reparameterizing $\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{1};\theta)$ as follows:
$\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{1};\theta)=\frac{\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)-\mathbf{x}_{t}}{1-t}$ (10)
Since both $\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)$ and $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$ are easy to sample, we use a single neural network with two separate output heads: the first head, $\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)$, to approximate the data samples, and the second, $\hat{\bm{\epsilon}}(\mathbf{x}_{t},t;\theta)$, to approximate the score function. The network generates these outputs simultaneously. The score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{1})$, using a Gaussian $p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{1})$, can be simplified to $-\frac{\bm{\epsilon_{0}}}{\sigma_{t}}$, where $\bm{\epsilon_{0}}\sim\mathcal{N}(\bm{0},\mathbf{\text{I}})$. The neural network can be trained using the following loss function:
@@
$\mathcal{L}_{\text{IDFF}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{x}_{t}}\big{[}\beta(t)^{2}|\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)-\mathbf{x}_{1}|^{2}+\lambda(t)^{2}\left|\hat{\bm{\epsilon}}(\mathbf{x}_{t},t;\theta)-\nabla_{x}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})\right|^{2}\big{]}$ (11)
@@
Where ${t\sim\mathcal{U}(0,1),\mathbf{x}_{0}\sim\mathcal{N}(\bm{0},\mathbf{\text{I}}),\mathbf{x}_{1}\sim p_{\text{data}},\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})}$. Here ${\lambda(\cdot),\beta(\cdot)}$ are two sets of positive weights.
Our proposed loss function comprises two terms: The first term optimizes the denoising model $\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t;\theta)$, with $\mathbf{x}_{t}$ as the noisy input. The second term represents a scoring model that approximates $\nabla_{x}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$. The first term of the loss incorporates a $\beta(t)$ weighting schedule, which emphasizes the quality of samples as they approach $t=1$. For more details about the loss function, please refer to Appendix A.1.2. The complete training process is described in Algorithm 1.
Sampling
@@
After training $\hat{\mathbf{x}}_{1}(.;\theta)$ and $\bm{\epsilon}(.;\theta)$, we can generate samples from IDFF by solving an SDE starting from the source samples $\mathbf{x}_{0}$. Refer to Figure 2.C for visualization. This procedure is outlined in Algorithm 2. The SDE is defined as $d\mathbf{x}_{t}=\mathbf{w}_{t}dt+\sigma_{t}d\mathbf{w}$, with initial samples $\mathbf{x}_{0}$ and target samples $\mathbf{x}_{1}$ at $t=1$. Therefore, we can draw samples from $p(\mathbf{x}_{0})$, run them through the SDE, and obtain the final samples $\mathbf{x}_{1}$. First, we draw $\mathbf{x}_{0}\sim\mathcal{N}(0,\mathbf{I})$, then compute $\hat{\mathbf{x}}_{1}(\mathbf{x}_{0},0,k;\theta)$. Next, we calculate $\hat{\mathbf{v}}_{t}$ using equation 10, followed by the computation of $\mathbf{w}_{t}$. After calculating $\mathbf{w}_{t}$ and using the SDE, we compute the mean for the next time step as $\bm{\mu}_{t+\Delta t}=(\mathbf{x}_{t}+\mathbf{w}_{t}\Delta t)$, and finally draw a sample from a small neighborhood of it: $\mathbf{x}_{t+\Delta t}\sim\mathcal{N}(\mathbf{x}_{t}+\mathbf{w}_{t}\Delta t,\sigma_{t}^{2}\Delta t\mathbf{I})$. The $\Delta t$ is determined by the NFE as $\Delta t=\frac{1}{\text{NFE}}$, where smaller NFE results in finer time discretization and tighter neighborhoods.
Likelihood calculation.
To evaluate the likelihood of data under IDFFs, we can leverage the change in the probability density as the sample ${\mathbf{x}}_{t}$ evolves according to the velocity field ${\mathbf{w}}_{t}$ as described by the continuity equation 4, which gives the time evolution of the probability density under the learned flow Lipman et al. (2022):
[…]
This integral is numerically approximated, often using Monte Carlo methods.
3.3 IDFF time-series adaptation
To use IDFF for time-series applications, we need to modify the training and sampling algorithms for IDFF as shown in Algorithms 4 and 3, respectively. To accomplish that, we define another random variable, $k$, which is a discrete variable and can take values of ${1,...,K}$ with equal probability, where $K$ is the length of a time series. Therefore, $k$ represents the index of the data sample in a time series. We also have $t$ as a continuous time variable interpolated between two subsequent values of $k$ (i.e., $k$ and $k-1$). Therefore we have $\forall k\in{1,...,K}\rightarrow p(t\mid k)\sim\mathcal{U}(0,1)$. This conditioning does not change the original assumption over $t$ since $k$ can take any values up to the time-series length with the same probability. With this in hand, we can directly pass the $k$ as additional inputs to $\hat{\bm{\epsilon}}_{1}(\mathbf{x}_{t},t,k;\theta)$ and $\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},t,k;\theta)$ and suggest Algorithm 1 for training and Algorithm 2 for sampling time-series data with IDFF. Note that for static data, the sequence length ($K$) equals one; otherwise, it equals the time series length.